solve the equation: log a (3x + 5) - log a (x - 5) = log a 8
Answer by Prithwis(166)

log a(3x + 5) - log a (x - 5) = log a 8
=> log a[(3x+5)/(x-5)] = log a 8
=> (3x+5)/(x-5) = 8
=> 3x+5 = 8x-40
=> 5x = 45
=> x = 9